OpenStudy (anonymous):

Integrate (2 ln x)/x

OpenStudy (anonymous):

do a u substitution. Let: \[u = \ln(x)\]

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{2lnx}{x}dx = ?\]

OpenStudy (anonymous):

put \(t = lnx\) then you have \(dx = dtx\)

OpenStudy (akshay_budhkar):

take lnx=u you will get dx=dt/x

OpenStudy (akshay_budhkar):

oh

OpenStudy (anonymous):

then the integral becomes \(2\int t dt\)\[2\frac{t^2}{2} + C\]

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{2\ln x}{x}dx = 2 \int\limits_{}^{}\ln x \frac{1}{x} dx \] substitute with\[u = \ln x\] how do we get \[\int\limits_{}^{}udu\] ?